OpenStudy (anonymous):

add and reduce: 2x+3/x^2-9+x/x-3

OpenStudy (anonymous):

3(x+1) /(x+4)(x-3) I believe this is right

OpenStudy (anonymous):

\[ \frac {(2x +3) + x(x+3)}{ (x^2 -9)}\] \[\frac {(2x+3+x^2 +3x)}{ (x^2 -9)}\] \[\frac {(x^2 +5x +3)} {(x^2 -9)}\]

OpenStudy (wwe123):